### v3

# Refinement Module based on Parse Graph of Feature Map for Human Pose Estimation

Abstract
Parse graphs of the human body can be obtained in the human brain to help humans complete the human Pose Estimation better (HPE). It contains a hierarchical structure, like a tree structure, and context relations among nodes. To equip models with such capabilities, many researchers predefine the parse graph of body structure to design HPE frameworks. However, these frameworks struggle to adapt to instances that deviate from the predefined parse graph and are often parameter-heavy. Unlike them, we view the feature map holistically, much like the human body. It can be optimized using parse graphs, where each node’s feature is an implicit expression rather than a fixed one. This allows it to adapt to more instances, unconstrained by rigid structural features. In this paper, we design the Refinement Module based on the Parse Graph of feature map (RMPG), which includes two stages: top-down decomposition and bottom-up combination. In the first stage, the feature map is decomposed into multiple sub-feature maps along the channel. In the second stage, the context relations of sub-feature maps are calculated to obtain their respective context information and the sub-feature maps with context information are concatenated along channels to obtain the refined feature map. Additionally, we design a hierarchical network with fewer parameters using multiple RMPG modules to model the context relations and hierarchies in the parse graph of body structure for HPE, some of which are supervised to obtain context relations among body parts. Our network achieves excellent results on multiple mainstream human pose datasets. More importantly, the effectiveness of RMPG is proven on different methods. The code of RMPG will be open.
Index Terms:
Parse Graph, refinement Module, top-down decomposition, bottom-up combination, context relations
I Introduction
The main task of 2D human posture estimation (HPE) is to obtain the positions of each joint of the human body in a image to determine the overall posture of the person. This is widely used in other tasks such as action recognition, character interaction, 3D human pose estimation, etc. We focus on single-person pose estimation.
Figure 1: (a) The parse graph of body structure, from [1]. The human body is partitioned into 5 parts (limbs and torso) and structured into 3 hierarchical levels (body, parts, joints), with semantic granularity decreasing from top to bottom. (b) The parse graph of feature maps.
When observing a person, humans can decompose the human body from the whole to part to primitives, thereby understanding the person more comprehensively. This structure can be expressed using parse graphs [2, 1]. As shown in Fig 1a, the parse graph includes the hierarchical structure and context relations among structures [2]. Hierarchical structure, as a tree-like decomposition, helps models capture multi-level relations, while context relations describe spatial interactions between parts to ensure spatial consistency. The parse graph is usually infered through top-down and bottom-up ways [2, 3, 4], which is also consistent with human visual patterns [5, 6]. In order for the model to have such capabilities, many methods [7, 8, 9, 10, 11, 12, 13, 14, 15, 16] try to use the context information or hierarchy in the parse graph of body structure for HPE. Liu et al. [1] think that the above methods do not model context relations and hierarchical structures simultaneously. So, they designed a new network to model context relations and hierarchies in the parse graph of body structure for HPE and achieve good results. However, its fixed body parse graph struggles with diverse human poses, and its more parameters limit its applicability. Different from them, we treat the feature map as a whole, like the human body. Then according to the parse graph of the feature map, the feature map is optimized through top-down decomposition and bottom-up combination (see Fig. 1b). Since the features of nodes can be learned implicit, which means the features are not fixed, it mitigates, to some extent, the problem of poor adaptability to various human poses.
In this paper, we design the Refinement Module based on the Parse Graph of feature map (RMPG) and use it to design a hierarchical network with fewer parameters than [1] to model the context relations and hierarchical in the parse graph of body structure (see Fig. 1a). For the RMPG, specifically, as shown in Fig. 1b, it consists of two stages, top-down decomposition and bottom-up combination. In the top-down decomposition stage, the feature map is decomposed along the channel dimension. The sub-feature maps also can be segmented which is an expansion along the depth direction. In the bottom-up combination stage, the context relations of sub-feature maps are calculated to obtain their respective context information and the sub-feature maps containing context information are concatenated to obtain the optimized feature map. Since it is optimizing feature maps, it can be easily applied by other methods. In addition, the RMPG module can help explicitly model the context relations and hierarchies in Fig. 1a. It only supervises sub-feature maps without the need to design additional network modules or frameworks. This is demonstrated by our hierarchical network, which consists of multiple RMPGs, some of which are supervised. In summary, the contributions of this paper are as follows:
  • We propose a new Refinement Module based on the Parse Graph of feature map (RMPG) to partially solve the problem that fixed body parse graphs are difficult to adapt to different samples. And the RMPG module effectiveness is demonstrated in other methods, including SimpleBaseline [17], Hourglass [18] and ViTPose [19].
  • We propose a hierarchical network based on RMPG to address the issue of excessive parameters when explicitly modeling contextual relationships and hierarchical structures in parse graphs.
  • Our hierarchical network is demonstrated on the CrowdPose, COCO keypoint detection, and MPII human pose datasets.
(a) The framework of our method.
(b) The base module of RMPG.
(c) Expand in depth direction of RMPG.
Figure 2: (a) Our network uses high level features to guide low level feature learning from the body to parts to joints, where $F_{1}$ comes from HRNet [20]. Through the RMPG module with supervision, the context relations among body parts can be obtained to help refine each body part (the yellow RMPG modules with supervision, and the blue RMPG modules without supervision). The heatmap within the red dashed box represents refined structural features from the previous RMPG module, with actual supervision information residing inside the RMPG module. Note: All heatmap supervision on feature maps is performed after dimensionality processing through a conv2d layer. (b) The RMPG module mainly operates on feature maps. First, the feature map is split into multiple sub-feature maps along the channel. Then the context relation of each child node is calculated to optimize each sub-node. The finally optimized child nodes are concatenated to obtain the refined feature map.
Figure 3: (a) Supervision from body to parts. (b) Supervision from parts to joints.
II Related work
Parse graphs. The parse graph includes a tree-structured decomposition in the vertical direction and context relations [2]. As shown in Fig. 1a, the human body is first divided into five parts, including limbs and torso. Furthermore, each part can continue to be decomposed. Many methods use hierarchical structures and context information in object parse graphs or other structure restraint to accomplish visual tasks, such as human pose estimation [21, 22, 23, 1, 15, 7, 8, 10, 11, 16], segmentation [24, 25, 26] and object detection [27, 24, 28]. Some of these methods [23, 24, 7, 27, 28, 25, 26, 15, 8, 10, 11, 16, 27] fail to simultaneously model the context relations and hierarchical structures in object parse graphs. Although PGBS [1] solves this problem with high parameters, it relies on fixed human structure splits and combinations, making it difficult to adapt to diverse postures. In addition, Chen et al. [22] use a discriminator to distinguish true and false postures, but training is difficult and Li et al. [21] use Transformer [29] to learn position relationship constraints, but its performance is limited on the MPII dataset [30], which may be because more data (e.g., the COCO keypoint detection dataset [31]) is required to achieve better results after introducing the Transformer. Different from these methods, we treat the feature map as an object and then optimize the feature map based on the parse graph. As shown in Fig. 1b, because it operates on feature maps, it can be easily used by other methods. In addition, as shown in Fig. 2a, RMPG also helps our network model context relations and hierarchies in the parse graph of body structure with lower parameters. Experiments demonstrate the effectiveness of our network and the generality of RMPG.
CNN based HPE. Some methods [20, 18, 32, 33, 17, 34, 35, 36] design many excellent network architectures for HPE. This method of improving backbone feature extraction capabilities for HPE seems to have become mainstream. However, these methods only rely on appearance features for learning, which can easily lead to network overfitting due to lacking the information of structure [2]. Inspired by [1], our hierarchical network as shown in Fig. 2a where $F_{1}$ comes from HRNet [20], we use multiple RMPG modules to model context relations and hierarchy in the parse graph of body structure.
Transformer based HPE. Transformer is proposed by [29] for neural machine translation (NMT) task [37]. It is widely used in various visual tasks and achieves good performance, such as HPE [19, 38, 21, 39]. Most methods focus on the application of transformer [39, 19, 40, 21] and simplifying transformer [41, 42]. On the contrary, our RMPG may provide possible optimization directions for the transformer. The transformer core is multi-headed attention and the heads are similar to nodes in a tree structure. However, compared with Fig. 1b, it lacks the context relations modeling for each head and does not further decomposition and combination for each head, which may be the direction of transformer [29] optimization. We add the RMPG module to the method of ViTPose [19]. Experimental results demonstrate the effectiveness of our module and also provide support for the possible optimization directions of the transformer structure mentioned above.
III Method
In this section, we first describe the derivation of the parse graph. Then, we present the overall network architecture. Subsequently, we detail the specific design of RMPG. Finally, we explain how to set up supervision in RMPG to capture context information between body parts.
III-A Parse graph
The parse graph includes hierarchical structures and contextual relations. It is represented as a 4-tuple $(V,E,\psi^{and},\psi^{leaf})$, where $(V,E)$ defines the graph structure, and $(\psi^{and},\psi^{leaf})$ are potential functions. Each node $u\in V$ has state variable $s_{u}=\{x_{u},y_{u}\}$, where $x_{u}$ is the position and $y_{u}$ is the type. The probability of the state variables $\Omega$ given an image $I$ is:
$P(\Omega|I)=\frac{1}{Z}\exp\{-E(\Omega,I)\}$ (1)
where $E(\Omega,I)$ is the energy function and $Z$ is the partition function. The energy function $F(\Omega)=-E(\Omega,I)$ is decomposed as:
$F(\Omega)=\sum_{u\in V_{L}}\psi^{leaf}_{u}(s_{u},I)+\sum_{u\in V_{A}}\psi^{and}_{u}(s_{u},{s_{v}}_{v\in C(u)})$ (2)
where $V_{L}$ and $V_{A}$ are leaf and non-leaf nodes respectively, and $C_{u}$ denotes the children of node $u$. The optimal state $\Omega^{*}$ is computed in two stages: bottom-up activation and top-down refinement. The bottom-up stage computes the maximum score $F_{u}^{\uparrow}(s_{u})$, while the top-down stage refines each node $v$ using its parent node $u$ and siblings:
$F_{v}^{\downarrow}(s_{v})=\psi_{u,v}(s_{u}^{*},s_{v})+\xi_{v}(s_{v},{s_{h}}_{h\in S_{v}})$ (3)
Where $S_{v}$ contains all nodes at the same level as $v$, $\xi_{v}$ captures context relations, and $s_{u}^{*}=\arg\max_{s_{u}}F_{u}^{\uparrow}(s_{u})$. This two-stage process ensures accurate part predictions by leveraging hierarchical and context information.
III-B Architecture
Our network explicitly models the context relations and hierarchical structure in Fig. 1a. As shown in Fig. 2a, our network is modified based on HRNet [20], where $F_{1}$ is the largest scale feature map in the fourth stage of HRNet. $F_{1}$ is used to generate the human body heatmap after a 2D convolution, which is supervised during training. We divide the human body into three levels, namely body, parts, and joints. For the parts and joints, we use the RMPG module with the supervision to obtain the context relations (see Fig. 3). After obtaining the refined feature map $F_{3}$ of the joint, we then use the unsupervised RMPG module to refine $F_{3}$ to obtain $F_{4}$. Finally, $F_{4}$ is used to generate the final joint heatmaps after a 2D convolution, which is supervised during training.
TABLE I: Comparisons on the COCO validation set.
Method Backbone #Params Input size MAP MAR
SimpleBaseline [17] ResNet-50 34.0M 256$\times$192 71.8 77.4
SimpleBaseline [17] ResNet-101 53.0M 256$\times$192 72.6 78.3
EMpose [43] HRNet-W32 30.3M 256$\times$192 75.0 80.2
OASNet [44] HRNet-W48 66.0M 256$\times$192 75.5 80.7
HRNet [20] HRNet-W32 28.5M 256$\times$192 74.4 79.8
HRNet [20] HRNet-W48 63.6M 256$\times$192 75.1 80.4
VITPose-B [19] ViT-B 86.0M 256$\times$192 75.8 81.1
TokenPose-L/D24 [21] HRNet-W48 27.5M 256$\times$192 75.8 80.9
PGBS [1] HRNet-W32 81.0M 256$\times$192 75.6 80.6
Ours-small HRNet-W32 37.1M 256$\times$192 75.3 80.4
Ours-large HRNet-W32 50.7M 256$\times$192 75.8 81.0
SimpleBaseline [17] ResNet-152 68.6M 384$\times$288 74.3 79.7
HRNet [20] HRNet-W32 28.5M 384$\times$288 75.8 81.0
HRNet [20] HRNet-W48 63.6M 384$\times$288 76.3 81.2
PGBS [1] HRNet-W32 81.0M 384$\times$288 76.9 81.7
Ours-small HRNet-W32 37.1M 384$\times$288 76.9 81.6
Ours-large HRNet-W32 50.7M 384$\times$288 77.4 82.1
Figure 4: (a) An input image of our network. (b) Visualization of context relations among all joint in (a). Each row represents the relation between a corresponding joint and all joints. The higher the score, the stronger the relation. From the visualization of the context relations in joints we notice that the joint tends to have a higher score with itself when it is visible , such as left hip, left wrist and left ankle. On the contrary, the joints tends to have a higher score with other joints when it is occluded or invisible, such as right wrist, right ankle and right shoulder.
III-C The design of RMPG
The RMPG module includes top-down decomposition and bottom-up composition, propagating context information from leaf nodes to higher levels to optimize feature maps, inheriting the hierarchical structure and context relations in Sec 3.1. The decomposition rule is to split the parent node features into individual child node features along the channel, which is a recursive operation. The combination rule is to recursively concatenate the child nodes with context information along the channel, which is a recursive operation.
TABLE II: Comparisons on the COCO test-dev set. $\dagger$ denotes the results of our reimplementation.
Method Backbone #Params Input size MAP MAR
TokenPose-L/D24 [21] HRNet-W48 27.5M 256$\times$192 75.1 80.2
EMpose [43] HRNet-W32 30.3M 256$\times$192 73.8 79.1
HRNet [20] HRNet-W32 28.5M 256$\times$192 73.5 78.9
ViTPose-B [19] ViT-B 86.0M 256$\times$192 75.1 78.3
PGBS [1] HRNet-W32 81.0M 256$\times$192 74.6 79.7
Ours-small HRNet-W32 37.1M 256$\times$192 74.4 79.5
Ours-big HRNet-W32 50.7M 256$\times$192 75.0 80.2
CPN (ensemble) [32] ResNet-Inception - 384$\times$288 73.0 79.0
SimpleBaseline [17] ResNet-152 68.6M 384$\times$288 73.7 79.0
TokenPose-L/D24 [21] HRNet-W48 29.8M 384$\times$288 75.9 80.8
[…]
HRNet [20] HRNet-W32 28.5M 384$\times$288 74.9 80.1
HRNet [20] HRNet-W48 63.6M 384$\times$288 75.5 80.5
PGBS [1] HRNet-W32 81M 384$\times$288 75.7 80.6
Ours-small HRNet-W32 37.1M 384$\times$288 75.8 80.7
Ours-large HRNet-W32 50.7M 384$\times$288 76.3 81.3
As shown in Fig. 2b, the feature map $F\in R^{C\times H\times W}$ is sliced into $N$ sub-feature maps $F_{1},F_{2},...,F_{N}\in R^{\frac{C}{n}\times H\times W}$ along the channel dimension, where $n=N=3$. Note that $n$ and $N$ can also be different (e.g., Eq. 8). In order to calculate the context relations among sub-feature maps in each sub-feature map, We reshape the $N$ sub-feature maps into $R^{L\times\frac{C}{n}}$, $L\in H\times W$, respectively. Then we concatenate the reshaped sub-feature maps to get $F_{all}\in R^{LN\times\frac{C}{n}}$. Finally the Context Information $CI()$ can be obtained:
$CI(\{F_{1},F_{2},\cdots,F_{N}\})=SoftMax(E)\times F_{all}$ (4)
where $E=F_{all}\times F_{all}^{T},E\in R^{LN\times LN}$ is the spatial correlation of the sub-feature maps, including the intra-correlation and inter-correlation and $\times$ represents matrix multiplication. $CI(\{F_{1},F_{2},\cdots,F_{N}\})\in R^{LN\times\frac{C}{n}}$ contains the context relations relations of each sub-features. Then, $CI(\{F_{1},F_{2},\cdots,F_{N}\})$ is reshaped in $R^{C\times H\times W}$. The refinement of the above RMPG module is completed. Finally, to preserve the properties of the original feature $F$, we add it as a residual to $CI$. The $CI()$ is $\xi_{v}=(s_{v},{s_{h}}_{h\in S_{v}})$ in Eq. 3.
The above is the refinement method when the depth is 1. When the depth is greater than 1, such as 2 (see Fig. 2c), the optimization process is as follows. First, there is the top-down decomposition process. The input feature map $F\in R^{C\times H\times W}$ will be decomposed to obtain all node features according to the setting of $gp=[3,3]$ (see Sec. 4.2):
$\begin{split}N=\{F,V_{1}^{1},V_{1}^{2},V_{1}^{3},V_{0}^{1},V_{0}^{2},\cdots\}\end{split}$ (5)
where $N$ is the set of features of all nodes in the parse graph of $F$:
$ch(V_{1}^{j})$ =$=\{V_{0}^{3(j-1)+k}\}_{k=1}^{3},\quad j=1,2,3$ (6)
$ch(F)$ =$=\{V_{1}^{1},V_{1}^{2},V_{1}^{3}\}$ (7)
where $ch$ represents child nodes. Child nodes are obtained by decomposing the parent node along the channel, so $V_{0}^{i}\in R^{\frac{C}{9}\times H\times W},i={1,2,\cdots,9}$ and $V_{1}^{m}\in R^{\frac{C}{3}\times H\times W},m={1,2,3}$. Then there is the bottom-up combination process. The context relations between leaf nodes belonging to the same parent node are calculated according to Eq. 4:
$C_{0}^{j}=CI(ch(V_{1}^{j})),\quad j=1,2,3$ (8)
where $C_{0}^{j}\in R^{L3\times\frac{C}{9}}$ is reshaped in $R^{\frac{C}{3}\times H\times W}$, which can be viewed as the node $V_{1}^{j}$ with context information, represented by $V_{1}^{j^{*}}$. Then the context information among $V_{1}^{j^{*}}$ is also calculated:
$F_{r}=CI(\{V_{1}^{j^{*}}\}_{j=1}^{3})$ (9)
where $F_{r}\in R^{L3\times\frac{C}{3}}$ is reshaped in $R^{C\times H\times W}$. The refinement of the above RMPG module is completed. Finally, to preserve the properties of the original feature $F$, we add it as a residual to $F_{r}$. The reasoning for other different $gp$ settings is similar to the above.
TABLE III: Comparisons on CrowdPose test set with YOLOv3 [45] human detector. * denotes using a stronger Faster RCNN [46] detector. $\dagger$ denotes the results of our reimplementation.
Method Backbone Input size MAP
SimpleBaseline [17] ResNet-152 256$\times$192 65.6
HRNet [20] HRNet-W32 256$\times$192 67.5
[…]
PGBS [1] HRNet-W32 256$\times$192 68.9
[…]
MIPNet [47] ResNet-101 384$\times$288 68.1
MIPNet* [47] ResNet-101 384$\times$288 70.0
HRNet* [20] HRNet-W48 384$\times$288 69.3
PGBS [1] HRNet-W32 384$\times$288 70.5
Ours-small HRNet-W32 256$\times$192 68.3
Ours-large HRNet-W32 256$\times$192 69.0
Ours-small HRNet-W32 384$\times$288 70.0
Ours-large HRNet-W32 384$\times$288 70.7
III-D Supervision
Supervision in our network. Firstly, we use the body heatmap to supervise the feature map $F_{1}$ after a 2D convolution (see Fig. 2a). Then, $F_{1}$ is decomposed (see Fig. 3a) according to the setting of $gp=[5,2]$:
$\begin{split}N=\{V_{n},V_{n-1}^{1},V_{n-1}^{2},\cdots\}\end{split}$ (10)
where $N$ is the set of nodes in the parse graph of $F_{1}$ and $n$ is the depth of the parse graph $n=len(gp)=2$. Next, perform supervision on all child nodes on the level 1:
$\begin{split}SV=\{V_{1}^{1},V_{1}^{2},V_{1}^{3},V_{1}^{4},V_{1}^{5}\}\end{split}$ (11)
where $V_{1}^{1}$ and $V_{1}^{2}$ are supervised by the left and right legs respectively, $V_{1}^{3}$ and $V_{1}^{4}$ are supervised by the left and right limbs respectively, and $V_{1}^{5}$ is supervised by torso. In this way, the context relations between parts can be obtained. Then, the refined feature maps $F_{2}$ (see Fig. 2a) are obtained, which is the result of the refinement containing all parts. Perform the same operation on $F_{2}$ as $F_{1}$ (see Fig. 3b), the difference is that the supervision information becomes joints. For example, $V_{1}^{1}$ is supervised by all joints on the left leg, including the left knee, left ankle, and left hip, and $V_{1}^{3}$ is supervised by all joints on the left limb, including the left shoulder, left elbow, and left wrist. Then, the refined feature maps $F_{3}$ (see Fig. 2a) are obtained, which is the result of the refinement containing all joints. Finally, $F_{3}$ is refined by the unsupervised RMPG module to obtain the final refinement feature maps $F_{4}$. Then $F_{4}$ goes through a 2d convolution to get the final joints prediction result.
Designing supervision. Follow the method of PGBS [1], the body heatmap is generated by placing a Gaussian kernel centered at the ground truth bounding box of the body and the size of the Gaussian kernel is proportional to the size of the human body in the image, and the parts heatmaps are generated by placing Gaussian kernels at the midpoints of skeletal segments, with kernel sizes proportional to bone lengths. For example, the left leg heatmap includes Gaussian kernels at the midpoints of the left hip-left knee and left knee-left ankle segments. During training, the score graphs $F_{v}^{\downarrow}(s_{v})$ in Eq. 3, which represent the heatmaps, are used as the ground truth to supervise our network.
IV Experiments
TABLE IV: Comparisons of PCKh@0.5 scores on the MPII test set. $\dagger$ denotes our replicated results.
Method Head Sho. Elb. Wri. Hip Knee Ank. Mean
De Bem et al. [15] 97.7 95.0 88.1 83.4 97.9 82.1 78.7 88.1
Luvizon et al. [48] 98.1 96.6 92.0 87.5 90.6 88.0 82.7 91.2
TokenPose-L/D6$\dagger$ [21] 98.4 96.3 91.7 87.2 90.5 87.7 83.5 91.1
Wang et al. [49] 98.3 96.7 92.4 88.5 90.4 88.3 84.4 91.6
Chou et al. [50] 98.2 96.8 92.2 88.0 91.3 89.1 84.9 91.8
Chen et al. [51] 98.1 96.5 92.5 88.5 90.2 89.6 86.0 91.9
Tang et al. [7] 98.4 96.9 92.6 88.7 91.8 89.4 86.2 92.3
HRNet-W32$\dagger$ 97.9 96.5 92.3 88.2 90.9 88.1 83.8 91.5
PGBS [1] 98.5 96.7 92.8 88.6 91.1 89.2 85.2 92.1
Ours-small 98.1 96.9 92.8 88.8 91.6 89.2 84.8 92.1
Ours-big 98.4 97.0 92.9 88.8 91.3 89.8 85.7 92.3
TABLE V: Ablation experiment comparison on the MPII test set. Set up unsupervised RMPG modules only. The supervised RMPG module settings remain unchanged, setting $gp=[5,2]$. $\dagger$ denotes the results of our reimplementation. $CR$ means context relations. Please note that multi-scale testing [52] is not used in ablation experiments.
Method Head Sho. Elb. Wri. Hip Knee Ank. Mean
Ours gp=[1] 98.4 96.7 92.2 87.9 90.5 88.8 84.4 91.6
Ours gp=[2] 98.4 96.5 92.3 87.7 90.5 88.5 84.5 91.6
Ours gp=[4] 98.3 96.5 92.4 88.1 90.5 89.0 84.4 91.7
Ours gp=[2,4] 98.4 96.6 92.3 88.2 90.9 89.1 84.6 91.8
Ours gp=[2,2] 98.5 96.7 92.5 88.2 91.0 89.0 84.6 91.9
Ours gp=[2,2] w/o $CR$ 98.3 96.6 92.0 87.9 90.5 88.1 83.8 91.4
Ours gp=[2,2,2] 98.6 96.5 92.1 88.0 90.6 88.2 84.3 91.6
HRNet-W32$\dagger$ [20] 98.3 96.4 91.9 87.7 90.2 87.7 83.7 91.2
TABLE VI: RMPG performance on different methods. The results are compared on the COCO validation set. The $gp$ represents the number of nodes and the depth settings in RMPG. The red upward arrow and numerical value indicate the improvement compared to the baseline network.
Method Backbone #Params Input size MAP MAR
SimpleBaseline [17] ResNet-50 34.0M 256$\times$192 71.8 77.4
SimpleBaseline [17] ResNet-101 53.0M 256$\times$192 72.8 78.3
Hourglass-52 [18] Hourglass-52 94.8M 256$\times$256 72.6 78.0
ViTPose [19] ViTPose-B 90.0M 256$\times$192 75.8 81.1
SimpleBaselines+RMPGgp=[2,2] ResNet-50 38.0M 256$\times$192 72.5 ($\uparrow$0.7) 78.0
SimpleBaselines+RMPGgp=[4,4] ResNet-50 36.7M 256$\times$192 72.5 ($\uparrow$0.7) 77.9
SimpleBaselines+RMPGgp=[2,4] ResNet-50 37.1M 256$\times$192 72.4 ($\uparrow$0.6) 77.9
SimpleBaselines+RMPGgp=[4,2] ResNet-50 36.7M 256$\times$192 72.6 ($\uparrow$0.8) 78.2
SimpleBaselines+RMPGgp=[2,2,2] ResNet-50 37.6M 256$\times$192 72.3 ($\uparrow$0.5) 77.9
SimpleBaselines+RMPGgp=[4,4,4] ResNet-50 36.8M 256$\times$192 72.3 ($\uparrow$0.5) 77.9
SimpleBaselines+RMPGgp=[2,2] ResNet-101 57.0M 256$\times$192 73.1 ($\uparrow$0.3) 78.7
SimpleBaselines+RMPGgp=[4,4] ResNet-101 55.7M 256$\times$192 72.8 ($\uparrow$0.0) 78.3
SimpleBaselines+RMPGgp=[4,2] ResNet-101 55.7M 256$\times$192 73.4 ($\uparrow$0.6) 78.8
SimpleBaselines+RMPGgp=[2,2,2] ResNet-101 56.6M 256$\times$192 72.8 ($\uparrow$0.0) 78.4
Hourglass+RMPGgp=[2,2] Hourglass-52 98.0M 256$\times$256 74.0 ($\uparrow$1.4) 79.4
Hourglass+RMPGgp=[4,4] Hourglass-52 97.6M 256$\times$256 73.8 ($\uparrow$1.2) 79.1
Hourglass+RMPGgp=[4,2] Hourglass-52 97.6M 256$\times$256 73.8 ($\uparrow$1.2) 79.2
Hourglass+RMPGgp=[2,2,2] Hourglass-52 98.4M 256$\times$256 73.8 ($\uparrow$1.2) 79.2
ViTPose+RMPGgp=[2,2] ViTPose-B 117.9M 256$\times$192 76.1 ($\uparrow$0.3) 81.3
ViTPose+RMPGgp=[4,4] ViTPose-B 114.0M 256$\times$192 76.0 ($\uparrow$0.2) 81.3
ViTPose+RMPGgp=[4,2] ViTPose-B 114.0M 256$\times$192 75.9 ($\uparrow$0.1) 81.2
ViTPose+RMPGgp=[2,2,2] ViTPose-B 121.8M 256$\times$192 76.0 ($\uparrow$0.2) 81.4
IV-A Datasets and evaluation methods
Datasets. The CrowdPose, COCO keypoint detection and MPII Human Pose datasets are trained and tested respectively in our method. For the CrowdPose datasets, there are $20$k images and $80$k human instances labeled with $14$ joints and the training, validation and testing subset are split in proportional to 5:1:4 [53]. For the COCO keypoint detection dataset, there are more than $200$k images and $250$k person instances, labeled with $17$ joints, of which $57$k images are used for training, $5$k images are used for validation, and $20$k images are used test. For the MPII Human Pose dataset, there are approximately $25$k images and $40$k annotated samples with $16$ joints per instance, of which $28$k are used for training and $11$k for testing.
Evaluation methods. For CrowdPose and COCO datasets, we use mean average precision (MAP) and mean average recall (MAR) when evaluating the model. In contrast, the MPII dataset uses PCKh score to evaluate the accuracy of pose estimation.
IV-B Implementation details
For the CrowdPose and COCO datasets, all input images are resized into $256\times 192$ or $384\times 288$ resolution. In verification and testing, we use YOLOv3 [45] human detector in the CrowdPose dataset and we use the detected person boxes [17] in the COCO dataset. For the MPII dataset, all input images are resized into $256\times 256$ resolution. In verification and testing, we use the provided person boxes and a six-scale pyramid testing method is used [52]. Other training and testing strategies are consistent with HRNet.
We use $gp$ to represent the settings of the tree structure in RMPG, including the number of nodes and depth. For example, $gp=[n_{1},n_{2},n_{3}]$ represents a tree with a depth of $gp$ list length $3$, in which there are $n_{1}$ nodes in the first level, each node in the first level has $n_{2}$ child nodes, and each node in the second level has $n_{3}$ child nodes. To be specific, Fig. 2b is $gp=[3]$ and Fig. 2c is $gp=[3,3]$. In our network, for supervised RMPG modules, setting $gp=[5,2]$, where 5 is the number of sub-feature maps after the first decomposition corresponding to the 5 body parts, and they are supervised by different body parts during training. For unsupervised RMPG modules, setting $gp=[2,2]$.
IV-C Benchmark results
Our network is divided into small model and large model. The main difference between them is that the number of convolution layers are different, resulting in changes in complexity, but there is no difference in the overall structure
COCO keypoint detection benchmark. Tables I and II show the results of our method and existing advanced methods on the coco validation and test-dev sets. For the COCO validation set, Our small model achieve $75.3$ MAP with the input size of $256\times 192$ and $76.9$ MAP with the input size of $384\times 288$, which are $0.9$ and $1.1$ higher than the baseline HRNet-W32 respectively. Our large model achieve $75.8$ MAP with the input size of $256\times 192$ and $77.4$ MAP with the input size of $384\times 288$, which are $1.4$ and $1.6$ higher than the baseline HRNet-W32 respectively, and $0.2$ and $0.5$ higher than the method of PGBS [1] respectively. For the COCO test-dev set, our small model achieve $74.4$ MAP with the input size of $256\times 192$ and $75.8$ MAP with the input size of $384\times 288$, which are $1.1$ and $0.9$ higher than the baseline HRNet-W32 respectively. Furthermore, our small model exceeds PGBS [1] by 0.1 when the input is $384\times 288$. Our large model achieves $75.0$ MAP with the input size of $256\times 192$ and $76.3$ MAP with the input size of $384\times 288$, which are $1.5$ and $1.4$ higher than the baseline HRNet-W32 respectively, and $0.4$ and $0.6$ higher than the method of PGBS [1] respectively. Compared with PGBS and ViT-B, our network can achieve better results with lower complexity.
CrowdPose benchmark. Table III shows the results of our method and existing advanced methods on the CrowdPose test set. Our small model achieves $68.3$ MAP with the input size of $256\times 192$ and $70.0$ MAP with the input size of $384\times 288$, which are $0.8$ and $0.7$ higher than the baseline HRNet respectively. Our large model, with low parameters, achieves $69.0$ MAP with the input size of $256\times 192$ and $70.7$ MAP with the input size of $384\times 288$, which are $0.1$ and $0.2$ higher than the method of PGBS [1] respectively.
MPII benchmark. As shown in Tabel IV, the PCKh@0.5 results of our method and other state-of-the-art methods on the MPII test set. By reproducing HRNet [20] with the input size of $256\times 256$, we achieve a $91.5$ PCKh@0.5 score with multi-scale testing [52]. Our small model achieves a $92.1$ PKCh@0.5 score with the input size of $256\times 256$, which is $0.6$ higher than the baseline HRNet and $1.0$ higher than TokenPose [21]. Our big model achieves a $92.3$ PKCh@0.5 score with the input size of $256\times 256$, which is $0.8$ higher than the baseline HRNet and $0.2$ higher than the method of [1].
Visualization of context relations. As shown in Fig. 4b, the visualization data of context relations from the supervised sub-feature map in Fig. 3b with the help Eq. 4, we can see that joints tend to have a higher context relations with itself when visible. Conversely, joints tend to have a higher context relations with other joints when they are occluded or invisible.
IV-D Ablation study
Our network ablation experiments are made on the MPII test set without multi-scale testing [52]. Mean PCKh@0.5 is used as the evaluation metric. In ablation experiments, our small model is used. All results are obtained with the input size of $256\times 256$. Tabel V shows the PCKh@0.5 results of our ablative experiment.
It can be seen from Tabel V that an appropriate increase in the number of leaves is beneficial for improving the results, such as the comparison between $gp=[1]$ and $gp=[4]$. An appropriate increase in depth is also beneficial for improving the results, such as $gp=[2,4]$ and $gp=[2,2]$. In addition, we also explore more about the impact of the number and depth of branches on the results in the next subsection.
In order to verify the effectiveness of the context relations, we also remove the context relations in the parse graph. When $gp=[2,2]$ without context relations, we get 91.4 PCKh@0.5, which is reduced by 0.5. This proves that the context relations is valid.
IV-E RMPG performance on different methods
To further demonstrate the efficacy of our proposed RMPG module, the RMPG module is used by us in other methods, including SimpleBaseline [17], Hourglass [18] and ViTPose [19].
The result of the RMPG on different methods as shown in Tabel VI, for the method of Hourglass, SimpleBaselines and ViTPose, the results are improved after adding the RMPG. To be specific, for the method of SimpleBaselines, we achieve 0.7 MAP improvement after introducing the RMPG module ($gp=[2,2]$ and $gp=[4,4]$) and 0.8 MAP improvement after introducing the RMPG module ($gp=[4,2]$) when the backbone is ResNet-50. What’s more, we achieve 0.3 MAP improvement after introducing the RMPG module ($gp=[2,2]$) and 0.6 MAP improvement after introducing the RMPG module ($gp=[4,2]$) when the backbone is ResNet-101. For the method of Hourglass [18], we achieve 1.4 MAP improvement after introducing the RMPG module ($gp=[2,2]$) and 1.2 MAP improvement after introducing the RMPG module ($gp=[4,4]$, $gp=[4,2]$ and $gp=[2,2,2]$) when the backbone is Hourglass-52. For the method of ViTPose [19], we achieve 0.3 MAP improvement after introducing the RMPG module ($gp=[2,2]$) when the backbone is ViTPose-B. The results of other $gp$ settings are shown in the Tabel VI. The above results prove the effectiveness of the RMPG module.
V Conclusion
The RMPG module based on parse graphs can obtain context relations between sub-feature maps, optimizing the network for better performance. In addition, RMPG can help to explicitly model context relations and hierarchies with lower parameters in parse graph of body structure. The experimental results demonstrate the effectiveness of our method, and we hope that the RMPG module can be widely used in various vision tasks.